# Diffusion approximation of the stationary distribution of a two-level single server queue

### v6

@@ 4.2 Preliminary lemmas
Thus, (4.65) is satisfied for $i=1,2$.
We next consider the case that $\lambda_{1}=\lambda_{2}$, equivalently, $c_{1}=c_{2}$. Then, from the definitions of $\Delta^{(r)}_{i}$, we have
$\Delta^{(r)}_{1}-\Delta^{(r)}_{2}=(\lambda^{(r)}_{1}-\lambda^{(r)}_{2})\mathbb{E}^{(r)}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]=o(r),$ (4.73)
because $\lim_{r\downarrow 0}\lambda^{(r)}_{1}-\lambda^{(r)}_{2}=0$ and (4.44). If $b_{1}=0$, then (4.54) and (4.60) yield
$\alpha^{(r)}_{e}(\Delta^{(r)}_{1}-\Delta^{(r)}_{2})=\frac{r\mu c_{1}\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)-r\mu b_{2}\varphi^{(r)}_{2}(0)+o(r),$
Hence, by (4.3) and $\varphi^{(r)}_{1}(0)+\varphi^{(r)}_{2}(0)=1$,
$\frac{\mu c_{1}\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)=\mu b_{2}(1-\varphi^{(r)}_{1}(0))+o(1)$
Since $c_{1}=c_{2}$, this proves (4.71). Similarly, if $b_{1}\not=0$, then (4.54), (4.63) and (4.3) yield
$\frac{\mu b_{1}}{e^{\beta_{1}\ell_{0}}-1}\varphi^{(r)}_{1}(0)-\mu b_{2}\varphi^{(r)}_{2}(0)=o(1).$
Again by $c_{1}=c_{2}$ and $\varphi^{(r)}_{1}(0)+\varphi^{(r)}_{2}(0)=1$, this proves (4.72).
Let
$\widetilde{\varphi}_{i}(\theta)=\widetilde{\varphi}_{i|cd}(\theta)\widetilde{\varphi}_{i}(0),\qquad i=1,2,$
then $\widetilde{\varphi}_{i}$ is the moment generating function of $\widetilde{\nu}_{i}$, so (4.68) and (4.69) together with (4.71) and (4.71) prove that $\nu^{(r)}_{i}$ weakly converges to $\widetilde{\nu}_{i}$ for $i=1,2$. Thus, $\nu^{(r)}$ weakly converges to $\widetilde{\nu}$ of (3.8). Similarly, by the assumption (3.c), (4.5), (4.59) and (4.62), we have
$\lim_{r\downarrow 0}\alpha^{(r)}_{e}=\lim_{r\downarrow 0}(c^{(r)}_{1}\mu^{(r)}(\varphi^{(r)}_{1}(0)-\mathbb{P}[L^{(r)}=0])+c^{(r)}_{2}\mu^{(r)}\varphi^{(r)}_{2}(0))=\lambda_{1}\widetilde{\varphi}_{1}(0)+\lambda_{2}\widetilde{\varphi}_{2}(0).$
@@
It remains to prove (iii). Assume that $\lambda_{1}>\lambda_{2}$, then $\Delta^{(r)}_{1}>\Delta^{(r)}_{2}$ by (4.73). For $b=0$, this yields
$\frac{\mu c_{1}\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)>\mu c_{2}b_{2}(1-\varphi^{(r)}_{1}(0))+o(1),$
@@
and therefore $\varphi^{(r)}_{1}(0)>\widetilde{A}_{1}>0$. Similarly, for $b\not=0$, we also have $\varphi^{(r)}_{1}(0)>\widetilde{A}_{1}>0$. Hence, (4.65) holds for $i=1$, so $\nu^{(r)}_{1|cd}$ weakly converges to $\widetilde{\nu}_{1|cd}$ as $r\downarrow 0$ by (i). Similarly, if $\lambda_{1}<\lambda_{2}$, then (4.65) holds for $i=2$, so $\nu^{(r)}_{2|cd}$ weakly converges to $\widetilde{\nu}_{2|cd}$.
@@ 5.1 Process limit
Up to now, we have considered the limit of the sequence of the scaled stationary distributions in heavy traffic. From the assumptions (3.a)–(3.e), one can guess this limiting distribution must be the stationary distribution of the reflected Brownian motion on $[0,\infty)$ whose Brownian component changes its drift and variance at the level $\ell_{0}$.
@@
$\sum_{i=1}^{2}\Big{(}\!-b_{i}\theta_{i}+\frac{1}{2}c_{i}\sigma_{i}^{2}\theta_{i}^{2}\Big{)}\widetilde{\varphi}_{i}(\theta_{i})+\left(\frac{\beta_{1}e^{\beta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}-\frac{\beta_{1}e^{\theta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}\right)\theta_{1}\frac{c_{1}\sigma^{2}_{1}}{2}\widetilde{\varphi}_{1}(0)$
@@
which follows from (4.3), (4.54) and (4.63) since $\beta_{i}=2b_{i}/(c_{i}\sigma_{i}^{2})$ for $i=1,2$. For $b_{1}=0$, BAR is obtained from (5.1) by letting $b_{1}\downarrow 0$. So, we only consider the case that $b_{1}\not=0$.
@@
where $W(\cdot)$ is the standard Brownian motion, $Y(\cdot)\equiv\{Y(t);t\geq 0\}$ is a non-deceasing process satisfying that $\int_{0}^{t}1(Z(u)>0)Y(du)=0$ for $t\geq 0$. If $Y(\cdot)$ is absolutely continuous with respect to Lebesgue measure on $\mathbb{R}_{+}$, then (5.2) has a unique weak solution by Theorem 5.15 of [6] because $\sigma(x)\geq\sqrt{c_{1}}\sigma_{1}\wedge\sqrt{c_{2}}\sigma_{2}>0$. However, $Y(\cdot)$ may not be absolutely continuous.
In the original submission of this paper, we assumed the existence of the solution $Z(\cdot)$ of (5.2), and considered its stationary distribution using the Ito formula, which yields
$\sum_{i=1}^{2}\Big{(}\!-b_{i}\theta+\frac{1}{2}c_{i}\sigma_{i}^{2}\theta^{2}\Big{)}\widehat{\varphi}_{i}(\theta)+\frac{\beta_{1}e^{\beta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}\theta\frac{c_{1}\sigma^{2}_{1}}{2}\widehat{\varphi}_{1}(0)=0,\qquad\theta\in\mathbb{R}.$ (5.3)
where $\widehat{\varphi}_{1}(\theta)=\mathbb{E}[e^{\theta Z}1(Z\leq\ell_{0})]$ and $\widehat{\varphi}_{2}(\theta)=\mathbb{E}[e^{\theta Z}1(Z>\ell_{0})]$ for random variable $Z$ subject to the stationary distribution of the $Z(\cdot)$. However, we had not solved this stationary equation. So, it was conjectured that
@@
    there is a unique weak solution $Z(\cdot)$ for the stochastic integral equation (5.2),
@@
    $L^{(r)}_{\rm dif}(\cdot)\equiv\{rL^{(r)}(r^{-2}t);t\geq 0\}$ weakly converges to the $Z(\cdot)$ of (a),
  • (c)
    the stationary distribution of $Z(t)$ uniquely exists under the stability condition $b_{2}>0$, and agrees with $\widetilde{\nu}$ of (3.8) in Theorem 3.1 under the one of conditions in (ii) of it in addition to $b_{2}>0$.
During the review process of this paper, the author have known that (a) is resolved by Atar et al. [1, Lemma 4.3] (see also Miyazawa [11] for a different proof), while (c) can be well answered because the stationary distribution of the $Z(\cdot)$ is fully obtained in [11], in which it is claimed that the limiting distribution $\widetilde{\nu}$ of this paper is identical with the stationary distribution of the $Z(\cdot)$, but this claim is only true if and only if $c_{1}=c_{2}$. This fact also can be seen from the comparison of (5.3) with (5.1), which does not require the $\widetilde{\nu}$ itself. However, the $\widetilde{\nu}$ has not yet been completely specified.
As for (a), the stationary distributions are obtained for more general reflected diffusion processes than $Z(\cdot)$. We refer to [11, 13] for them, and will not further discuss about the conjectures (a) and (c). As for the conjecture (b), it is still open.
@@
We only verify (3.1) for $b_{1}\not=0$ because it is similarly proved for $b_{1}=0$. First, from (4.62) and (3.10),
$\frac{1}{r}\mathbb{P}[L^{(r)}=0)]$ $=\frac{b_{1}e^{\beta_{1}\ell_{0}}}{c_{1}(e^{\beta_{1}\ell_{0}}-1)}\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}+o(1)$
@@
We first prove (4.22) and (4.4). Apply $f({\boldsymbol{x}})=x_{2}^{k+1}1(x_{2}\in B)$ to (4.1), then
$(k+1)\mathbb{E}[(R^{(r)}_{e})^{k}1(R^{(r)}_{e}\in B)]$ $=\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{e,1})^{k+1})1(T^{(r)}_{e,1}\in B)]\mathbb{P}^{(r)}_{e}[L^{(r)}(0)\leq\ell^{(r)}_{0}]$
$\quad+\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{e,2})^{k+1})1(T^{(r)}_{e,2}\in B)]\mathbb{P}^{(r)}_{e}[L^{(r)}(0)>\ell^{(r)}_{0}],$
where (4.1) is still applicable since the right-hand side of this equation is finite. Hence, (4.22) is obtained. (4.4) is similarly proved. We next prove (4.24).
@@
Hence, the last inequality in (A.3) for $k=1$ and the uniform integrability condition yield

### v7

@@ 4.3 Expansion of the asymptotic BAR
We next consider the case that $\lambda^{(r)}_{1}-\lambda^{(r)}_{2}=O(r)$. Define $\overline{\varphi}_{1}(0)$ and $\underline{\varphi}_{1}(0)$ as
$\overline{\varphi}_{1}(0)=\limsup_{r\downarrow 0}\varphi^{(r)}_{1}(0),\qquad\underline{\varphi}_{1}(0)=\liminf_{r\downarrow 0}\varphi^{(r)}_{1}(0),$
and let $\{\overline{r}_{n}\in(0,1];n\geq 1\}$ and $\{\underline{r}_{n}\in(0,1];n\geq 1\}$ be the sequences such that $\varphi^{(\overline{r}_{n})}_{1}(0)$ and $\varphi^{(\underline{r}_{n})}_{1}(0)$ attain $\overline{\varphi}_{1}(0)$ and $\underline{\varphi}_{1}(0)$ as $n\to\infty$, respectively. We aim to show $\overline{\varphi}_{1}(0)=\underline{\varphi}_{1}(0)$ using $\lambda^{(r)}_{1}-\lambda^{(r)}_{2}=O(r)$.
To see this, we first consider $\overline{\varphi}_{1}(0)$. Since $\varphi^{(\overline{r}_{n})}_{1}(0)$ converges to $\overline{\varphi}_{1}(0)$, applying (5.1) and (4.17) to Lemma 4.11, we have
$\lim_{n\to\infty}\varphi^{(\overline{r}_{n})}_{1}(\theta_{1})=\widetilde{\varphi}_{1}(\theta_{1})\lim_{n\to\infty}\varphi^{(\overline{r}_{n})}_{1}(0)=\widetilde{\varphi}_{1}(\theta_{1})\overline{\varphi}_{1}(0).$ (5.16)
Since the distribution $\widetilde{\nu}_{1}$ has the bounded density $h_{1|cd}$ on $[0,\ell_{0}]$, this implies that
$\lim_{n\to\infty}\mathbb{P}[L^{(\overline{r}_{n})}=\ell^{(\overline{r}_{n})}_{0}]\leq\lim_{n\to\infty}\mathbb{P}[\overline{r}_{n}L^{(\overline{r}_{n})}\in[\ell_{0}-\varepsilon,\overline{r}_{n}\ell^{(\overline{r}_{n})}_{0}]]$
[…]
because $0\leq\overline{\varphi}_{1}(0)\leq 1$. Hence, $\mathbb{P}[L^{(\overline{r}_{n})}=\ell^{(\overline{r}_{n})}_{0}]=o(1)$ as $n\to\infty$. By Lemma 5.1, this implies that
$\lim_{n\to\infty}\mathbb{E}^{(\overline{r}_{n})}_{s}[\widehat{R}^{(\overline{r}_{n})}_{e}1(L^{(\overline{r}_{n})}(0)=\ell^{(\overline{r}_{n})}_{0})]=0.$ (5.18)
Then, from the definitions of $\Delta^{(r)}_{i}$ and $\lambda^{(\overline{r}_{n})}_{1}-\lambda^{(\overline{r}_{n})}_{2}=O(\overline{r}_{n})$, we have
$\Delta^{(\overline{r}_{n})}_{1}-\Delta^{(\overline{r}_{n})}_{2}=(\lambda^{(\overline{r}_{n})}_{1}-\lambda^{(\overline{r}_{n})}_{2})\mathbb{E}^{(\overline{r}_{n})}_{s}[\widehat{R}^{(\overline{r}_{n})}_{e}1(L^{(\overline{r}_{n})}(0)=\ell^{(\overline{r}_{n})}_{0})]=o(\overline{r}_{n}),$ (5.19)
If $b_{1}=0$, then (4.50) and Lemma 4.10 yield
$\alpha^{(\overline{r}_{n})}_{e}(\Delta^{(\overline{r}_{n})}_{1}-\Delta^{(\overline{r}_{n})}_{2})=\frac{rc_{1}\mu\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(\overline{r}_{n})}_{1}(0)-r\mu b_{2}\varphi^{(\overline{r}_{n})}_{2}(0)+o(\overline{r}_{n}),$
Hence, by (5.19) and $\varphi^{(\overline{r}_{n})}_{1}(0)+\varphi^{(\overline{r}_{n})}_{2}(0)=1$,
$\frac{c_{1}\mu\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(\overline{r}_{n})}_{1}(0)=\mu b_{2}(1-\varphi^{(\overline{r}_{n})}_{1}(0))+o(1).$
This is identical with (5.12) for $r=\overline{r}_{n}$ because $c_{1}=c_{2}$ by $\lambda_{1}=\lambda_{2}$. Hence, we have (5.13) for $r=\overline{r}_{n}$ and $b_{1}=0$. Similarly, if $b_{1}\not=0$, then, again by (4.50), (5.11) and (5.19) for $r=\overline{r}_{n}$, we have
$\frac{b_{1}\mu}{e^{\beta_{1}\ell_{0}}-1}\varphi^{(\overline{r}_{n})}_{1}(0)-\mu b_{2}\varphi^{(\overline{r}_{n})}_{2}(0)=o(1).$
This is identical with (5.14) for $r=\overline{r}_{n}$ by $\lambda_{1}=\lambda_{2}$. Hence, we have (5.15) for $r=\overline{r}_{n}$ and $b_{1}\not=0$. Thus, we have
$\overline{\varphi}_{1}(0)=\lim_{n\to\infty}\varphi^{(\overline{r}_{n})}_{1}(0)=\widetilde{\varphi}_{1}(0)>0.$
We can repeat the exactly same arguments for $\{\underline{r}_{n};n\geq 0\}$, and therefore we also have
$\underline{\varphi}_{1}(0)=\lim_{n\to\infty}\varphi^{(\underline{r}_{n})}_{1}(0)=\widetilde{\varphi}_{1}(0)>0.$
Hence, $\overline{\varphi}_{1}(0)=\underline{\varphi}_{1}(0)$ as claimed, and therefore $\lim_{r\downarrow 0}\varphi^{(r)}_{1}(0)=\widetilde{\varphi}_{1}(0)>0$. Hence, the condition (3.6) in the part (i)’ is satisfied, so we can use the results in that part. Namely, we have (5.7) and (5.8), equivalently, (3.7). From this together with (5.13) and (5.15), we have (3.11).
It remains to prove (3.12) in the part (ii)’. To see this, apply the assumption (3.c), (5.13), (5.15) and Lemma 4.9 to (4.5), then we have
$\lim_{r\downarrow 0}\alpha^{(r)}_{e}=\lim_{r\downarrow 0}(c^{(r)}_{1}\mu^{(r)}(\varphi^{(r)}_{1}(0)-\mathbb{P}[L^{(r)}=0])+c^{(r)}_{2}\mu^{(r)}\varphi^{(r)}_{2}(0))=\lambda_{1}\widetilde{\varphi}_{1}(0)+\lambda_{2}\widetilde{\varphi}_{2}(0),$
which proves (3.12).
@@
Finally, we prove the part (iii)’. Assume that $\lambda_{1}>\lambda_{2}$, then $\Delta^{(r)}_{1}>\Delta^{(r)}_{2}$ by (5.19). For $b=0$, this yields
$\frac{c_{1}\mu\sigma_{1}^{2}}{2\ell_{0}}\varphi^{(r)}_{1}(0)>\mu c_{2}b_{2}(1-\varphi^{(r)}_{1}(0))+o(1),$
@@
and therefore $\varphi^{(r)}_{1}(0)>\widetilde{A}_{1}>0$. Similarly, for $b\not=0$, we also have $\varphi^{(r)}_{1}(0)>\widetilde{A}_{1}>0$. Hence, (3.6) holds for $i=1$, so $\nu^{(r)}_{1|cd}$ weakly converges to $\widetilde{\nu}_{1|cd}$ as $r\downarrow 0$ by the part (i)’. Similarly, if $\lambda_{1}<\lambda_{2}$, then (3.6) holds for $i=2$, so $\nu^{(r)}_{2|cd}$ weakly converges to $\widetilde{\nu}_{2|cd}$. Thus, the proof of Proposition 5.1 is completed.
5.2 Proof of Proposition 5.2
Using (5.2), we prove (5.1) by reductio ad absurdum. First, assume that ${\mathcal{E}}^{(r)}_{1}$ does not vanish as $r\downarrow 0$. Then, there are a sequence $\{r_{n}\in(0,1];n\geq 0\}$ and a constant $\delta_{1}\not=0$ such that $r_{n}\downarrow 0$ and ${\mathcal{E}}^{(r_{n})}_{1}\to\delta_{1}$ as $n\to\infty$, because ${\mathcal{E}}^{(r)}_{1}$ is uniformly bounded for $r\in(0,1]$ by Corollary 4.4.
We next choose a subsequence $\{r^{\prime}_{n}\in(0,1];n\geq 0\}\subset\{r_{n}\in(0,1];n\geq 0\}$ such that $\alpha^{(r)}_{e}$ and $\varphi^{(r^{\prime}_{n})}_{1}(0)$ converge to constants $a_{e}>0$ and $d_{1}\in[0,1]$, respectively, as $n\to\infty$, which are possible because $\alpha^{(r)}_{e}$ is uniformly bounded away from 00 and above by Corollary 4.1 and $\varphi^{(r^{\prime}_{n})}_{1}(0)\in[0,1]$. Then, by Lemma 4.11, we have
$\lim_{n\to\infty}\varphi^{(r)}_{1}(\theta_{1})$ =$\begin{cases}\frac{e^{\theta_{1}\ell_{0}}-1}{\ell_{0}\theta_{1}}\Big{(}d_{1}+\frac{2a_{e}\delta_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2a_{e}\delta_{1}}{c_{1}\mu\sigma_{1}^{2}}e^{\theta_{1}\ell_{0}},&b_{1}=0,\\ \frac{\beta_{1}}{e^{\beta_{1}\ell_{0}}-1}\frac{e^{\beta_{1}\ell_{0}}-e^{\theta_{1}\ell_{0}}}{\beta_{1}-\theta_{1}}\Big{(}d_{1}+\frac{2a_{e}\delta_{1}}{c_{1}\mu\sigma_{1}^{2}}\Big{)}-\frac{2a_{e}\delta_{1}}{c_{1}\mu\sigma_{1}^{2}}e^{\theta_{1}\ell_{0}},&b_{1}\not=0.\end{cases}$
This shows that $\nu^{(r)}_{1}$ weakly converges to a finite measure on $[0,\ell_{0}]$, but this measure cannot have a negative mass at $\ell_{0}$, so we must have that $\delta_{1}<0$ because $\delta_{1}\not=0$.
We next consider the sequence $\{{\mathcal{E}}^{(r^{\prime}_{n})}_{2};n\geq 1\}$. If this sequence does not vanish as $n\to\infty$, then there is a subsequence $\{r^{\prime\prime}_{n}\in(0,1];n\geq 0\}\subset\{r^{\prime}_{n}\in(0,1];n\geq 0\}$ such that ${\mathcal{E}}^{(r^{\prime\prime}_{n})}_{2}$ converges some $\delta_{2}\not=0$ as $n\to\infty$. Then, by the exactly same argument as used for $\{{\mathcal{E}}^{(r_{n})}_{1};n\geq 1\}$ using (4.57) instead of (4.56), we must have that $\delta_{2}<0$. Hence, we have
$\lim_{r^{\prime\prime}_{n}\downarrow 0}({\mathcal{E}}^{(r^{\prime\prime}_{n})}_{1}+{\mathcal{E}}^{(r^{\prime\prime}_{n})}_{2})=\delta_{1}+\delta_{2}<0.$
This contradicts (5.2), and therefore ${\mathcal{E}}^{(r)}_{1}$ must vanish as $r\downarrow 0$. By (5.2), this further implies that ${\mathcal{E}}^{(r)}_{2}$ vanishes as $r\downarrow 0$. Hence, we have (5.1).
We now prove that (3.f) implies (5.2). By the definitions of ${\mathcal{E}}^{(r)}_{1}$ and ${\mathcal{E}}^{(r)}_{2}$, we have
${\mathcal{E}}^{(r)}_{1}+{\mathcal{E}}^{(r)}_{2}$ $=-2^{-1}\mathbb{E}^{(r)}_{s}[\{((\lambda^{(r)}_{1}\sigma^{(r)}_{e,1})^{2}+2)(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{1}\widehat{R}^{(r)}_{e})^{2}\}1(L^{(r)}(0)=\ell^{(r)}_{0})]$
$\quad+2^{-1}\mathbb{E}^{(r)}_{s}[\{(\lambda^{(r)}_{2}(\sigma^{(r)}_{e,2})^{2}+2)(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})-(\lambda^{(r)}_{2}\widehat{R}^{(r)}_{e})^{2}\}1(L^{(r)}(0)=\ell^{(r)}_{0})].$ (5.20)
We separately consider the two cases that $\lambda_{1}=\lambda_{2}$ and $\lambda_{1}\not=\lambda_{2}$. First, consider the case that $\lambda_{1}=\lambda_{2}$. In this case, $\sigma_{e,1}^{2}=\sigma_{e,2}^{2}$ by (3.f). Hence, (5.2) implies (5.2).
Next consider the case that $\lambda_{1}\not=\lambda_{2}$. In this case, we consider the BAR (4.3) for $\theta_{2}=\theta_{1}=\theta\leq 0$. From (4.2), we have
$D^{(r)}(\theta,\theta)=\alpha^{(r)}_{e}e^{r\theta(\ell^{(r)}_{0}+1)}\mathbb{E}^{(r)}_{s}[(e^{-\eta^{(r)}_{1}(r\theta)\widehat{R}^{(r)}_{e}}-e^{-\eta^{(r)}_{2}(r\theta)\widehat{R}^{(r)}_{e}})1(L^{(r)}(0)=\ell^{(r)}_{0})].$ (5.21)
Applying (4.37), (4.8) and Corollary 4.4 to this equation, we have
$D^{(r)}(\theta,\theta)$ $=\alpha^{(r)}_{e}e^{\theta\ell_{0}}(1+r\theta+o(r\theta))\mathbb{E}^{(r)}_{s}\big{[}(r\theta(\lambda^{(r)}_{2}-\lambda^{(r)}_{1})\widehat{R}^{(r)}_{e}$
$\quad+2^{-1}r^{2}\theta^{2}\{(\lambda^{(r)}_{2}\sigma^{(r)}_{e,2})^{2}\lambda^{(r)}_{2}-(\lambda^{(r)}_{1}\sigma^{(r)}_{e,1})^{2}\lambda^{(r)}_{1}\}\widehat{R}^{(r)}_{e}$
$\quad-2^{-1}r^{2}\theta^{2}\{(\lambda^{(r)}_{2})^{2}-(\lambda^{(r)}_{1})^{2}\}(\widehat{R}^{(r)}_{e})^{2})1(L^{(r)}(0)=\ell^{(r)}_{0})\big{]}+o((r\theta)^{2}),$
$=\alpha^{(r)}_{e}e^{\theta\ell_{0}}\big{\{}r\theta(\lambda^{(r)}_{2}-\lambda^{(r)}_{1})\mathbb{E}^{(r)}_{s}\big{[}\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})\big{]}+(r\theta)^{2}({\mathcal{E}}^{(r)}_{1}+{\mathcal{E}}^{(r)}_{2})\big{\}}+o((r\theta)^{2}).$
This implies that
$D^{(r)}(\theta,\theta)$ $=r\theta\alpha^{(r)}_{e}e^{\theta\ell_{0}}(\lambda^{(r)}_{2}-\lambda^{(r)}_{1})\mathbb{E}^{(r)}_{s}\big{[}\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})\big{]}+O((r\theta)^{2}).$ (5.22)
Substitute this formula into the BAR (4.3) for $\theta_{2}=\theta_{1}=\theta\leq 0$, which is given by
$\frac{1}{2}\sum_{i=1,2}c_{i}\sigma_{i}^{2}\big{(}(\beta_{i}-\theta)\mu\theta r^{2}+o((\theta r)^{2})\big{)}\psi^{(r)}_{i}(\theta)$
$\quad=c_{1}\left(\mu r\theta-\mu^{3}\sigma_{s}^{2}r^{2}\theta^{2}/2+o((\theta r)^{2})\right)\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]+D^{(r)}(\theta,\theta).$ (5.23)
Then, picking up the terms of order $r\theta$, we have
$c_{1}\mu\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]+\alpha^{(r)}_{e}e^{\theta\ell_{0}}(\lambda_{2}-\lambda_{1})\mathbb{E}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]=O(r\theta).$
Since $\mathbb{E}[1(L^{(r)}=0)g^{(r)}_{1,r\theta}(R^{(r)})]=O(r)$ by Corollary 4.2 and Lemma 4.9, this yields
$(\lambda_{2}-\lambda_{1})\mathbb{E}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]=O(r).$ (5.24)
Since $\lambda_{1}\not=\lambda_{2}$ by (f.1), (5.24) implies (5.4), which further implies (5.5). Hence,
$\lim_{r\downarrow 0}\mathbb{E}_{s}[\widehat{R}^{(r)}_{e}1(L^{(r)}(0)=\ell^{(r)}_{0})]=0,$ (5.25)
and therefore we have (5.2) by (5.2) and Lemma 5.2. Thus, the proof of Proposition 5.2 is completed.
@@ 6.1 Process limit
Up to now, we have considered the limit of the sequence of the scaled stationary distributions in heavy traffic. From the assumptions (3.a)–(3.f), one can guess this limiting distribution must be the stationary distribution of the reflected Brownian motion on $[0,\infty)$ whose Brownian component changes its drift and variance at the level $\ell_{0}$.
@@
$\sum_{i=1}^{2}\Big{(}\!-b_{i}\theta_{i}+\frac{1}{2}c_{i}\sigma_{i}^{2}\theta_{i}^{2}\Big{)}\widetilde{\varphi}_{i}(\theta_{i})+\frac{\beta_{1}e^{\beta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}\theta_{1}\frac{c_{1}\sigma^{2}_{1}}{2}\widetilde{\varphi}_{1}(0)-\frac{\beta_{1}e^{\theta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}\theta_{1}\frac{c_{1}\sigma^{2}_{1}}{2}\widetilde{\varphi}_{1}(0)$
@@
which can be obtained from (4.3) and (4.2) similarly to $D^{(r)}(\theta,\theta)$ from (5.22) and (5.2), recall that $\beta_{i}=2b_{i}/(c_{i}\sigma_{i}^{2})$ for $i=1,2$. The BAR for $b_{1}=0$ is obtained from (6.1) by letting $b_{1}\downarrow 0$. So, we only consider the case that $b_{1}\not=0$.
@@
where $W(\cdot)$ is the standard Brownian motion, $Y(\cdot)\equiv\{Y(t);t\geq 0\}$ is a non-deceasing process satisfying that $\int_{0}^{t}1(Z(u)>0)Y(du)=0$ for $t\geq 0$. Recently, Miyazawa [11] (see also [1, 2, 13]) shows that (6.2) has a unique weak solution, and $Z(t)$ has a stationary distribution if only if $b_{2}>0$, which corresponds to $\rho^{(r)}_{2}<1$.
Assume that $Z(\cdot)$ is stationary, then, applying the Ito formula to (6.2) for test function $f(x)=e^{\theta x}$, we have
$\sum_{i=1}^{2}\Big{(}\!-b_{i}\theta+\frac{1}{2}c_{i}\sigma_{i}^{2}\theta^{2}\Big{)}\widehat{\varphi}_{i}(\theta)+\frac{\beta_{1}e^{\beta_{1}\ell_{0}}}{e^{\beta_{1}\ell_{0}}-1}\theta\frac{c_{1}\sigma^{2}_{1}}{2}\widehat{\varphi}_{1}(0)=0,\qquad\theta\leq 0.$ (6.3)
where $\widehat{\varphi}_{1}(\theta)=\mathbb{E}[e^{\theta Z}1(Z\leq\ell_{0})]$ and $\widehat{\varphi}_{2}(\theta)=\mathbb{E}[e^{\theta Z}1(Z>\ell_{0})]$ for random variable $Z$ subject to the stationary distribution of the $Z(\cdot)$. (6.3) is the stationary equation for $Z(\cdot)$. Then, we can see that (6.1) for $\theta_{1}=\theta_{2}=\theta$ agrees with (6.3) if and only if $c_{1}=c_{2}$, equivalently, $\lambda_{1}=\lambda_{2}$. Hence, if $L^{(r)}_{\rm dif}(\cdot)$ weakly converges to $Z(\cdot)$ as $r\downarrow 0$, then we must have $\lambda_{1}=\lambda_{2}$. Of course, this does not exclude the possibility that $L^{(r)}_{\rm dif}(\cdot)$ weakly converges to some process other than $Z(\cdot)$. We conjecture that this is indeed possible because the stationary distribution of $L^{(r)}_{\rm dif}(\cdot)$ weakly converges to the $\widetilde{\nu}$ of (3.9) for $\lambda_{1}\not=\lambda_{2}$ at least when $T_{s}$ is exponentially distribution by (ii) of Theorem 3.1. Thus, we have two challenging problems.
@@
    Prove that $L^{(r)}_{\rm dif}(\cdot)$ weakly converges to $Z(\cdot)$ when $\lambda_{1}=\lambda_{2}$.
@@
    Find a process limit of $L^{(r)}_{\rm dif}(\cdot)$ when $\lambda_{1}\not=\lambda_{2}$.
Both problems require theoretical investigations based on process limits, which is beyond the present framework. So, we leave them for future study.
@@
We only verify (3.1) for $b_{1}\not=0$ because it is similarly proved for $b_{1}=0$. Since ${\mathcal{E}}^{(r)}_{1}=o(1)$ by Proposition 5.2, it follows from (4.53) and (3.10) that
$\frac{1}{r}\mathbb{P}[L^{(r)}=0]$ $=\frac{b_{1}e^{\beta_{1}\ell_{0}}}{c_{1}(e^{\beta_{1}\ell_{0}}-1)}\frac{c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}{c_{2}b_{1}+c_{1}b_{2}(e^{\beta_{1}\ell_{0}}-1)}+o(1)$
@@
We first prove (4.26) and (4.27). Apply $f(z,{\boldsymbol{x}})=x_{1}^{k+1}$ to (4.1), then
$(k+1)\mathbb{E}[(R^{(r)}_{e})^{k}]$ $=\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{e,1})^{k+1})]\mathbb{P}^{(r)}_{e}[L^{(r)}(0)\leq\ell^{(r)}_{0}]$
$\quad+\alpha^{(r)}_{e}\mathbb{E}[(T^{(r)}_{e,2})^{k+1})]\mathbb{P}^{(r)}_{e}[L^{(r)}(0)>\ell^{(r)}_{0}],$
where (4.1) is applicable since the right-hand side of this equation is finite. Hence, (4.26) is obtained. Similarly, we have (4.27) applying $f(z,{\boldsymbol{x}})=x_{2}^{k+1}$ to (4.1) because
$f^{\prime}(X^{(r)}(t))=-(k+1)(R^{(r)}_{s}(t))^{k}(1-1(L^{(r)}(t)=0)),$
and $R^{(r)}(t)$ has the same distribution as $T^{(r)}_{s}$ when $L^{(r)}(t)=0$. We next prove (4.28). For this, we use the following inequality,
@@
Hence, the last inequality in (A.3) for $k=1$ and the uniform boundedness of $\mathbb{E}[(T^{(r)}_{u})^{3}$ yield